OpenStudy (anonymous):

What is the quadratic function that is created with roots at 2 and 4 and a vertex at (3, 1)?

OpenStudy (anonymous):

say the quadratic is, a(x-2)(x-4)

OpenStudy (anonymous):

y = a(x-2)(x-4)

OpenStudy (anonymous):

since it passes through vertex (3,1) plugin x=3, y=1 1 = a(3-2)(3-4) 1 = a(1)(-1) a = -1

OpenStudy (anonymous):

so the quadric wud be y = -1(x-2)(x-4) multiply and simplify

OpenStudy (anonymous):

y = -1(x-2)(x-4)? = -x2 + 6x − 8 (x) = -x2 + 6x − 8

OpenStudy (anonymous):

@sara12345